A gas occupies 200 cm³ at a temperature 27degree Celsius and 750 mm of Hg pressure.  Find the volume of gas at -3degree Celsius and 760 mm [Hg] pressure. 

Grade:9

1 Answers

Khimraj
5 years ago
using PV = nRT
P1*V1/P2*V2 = T1/T2
T1 = 27+273 = 300K
T2 = -3+273 = 270K
(750*200)/(760*V2) = 300/270
V2 = (750*200*270)/(760*300)
V2 = 177.63 cm3
Hope it clears...............................